OpenStudy (anonymous):

how many photons are produced in a laser pulse of0.472J at 633nm

OpenStudy (anonymous):

\[E _{photon}=\frac{ hc }{ \lambda }\] \[N _{photon}=\frac{ E }{ E _{photon} }\]

OpenStudy (anonymous):

h is Planck's constant, which is 6.626x10^-34 c is speed of light, which is 3x10^8 \lambda is the wavelength, which is 633nm E is the energy of the laser pulse

OpenStudy (anonymous):

\[E _{photon}=\frac{ hc }{ \lambda }=\frac{ 6.623\times 10^{-34}\times 3\times 10^{8} }{ 633\times 10^{-9} }\approx 3.14\times 10^{-19}(J)\] \[\rightarrow N _{photon}=\frac{ E }{ E _{photon} }=\frac{ 0.472 }{ 3.14\times 10^{-19} }\approx 1.50\times 10 ^{18}(photon)\]